# Text-to-Image Diffusion Models are Great Sketch-Photo Matchmakers

### 12 Mar 2024 v1

@@ 3 Revisiting Text-to-Image Diffusion Models
Stable Diffusion Architecture  Given an image and text-prompt pair ($\mathbf{x,p}$), SD [69] first uses the encoder $\mathcal{E}(\cdot)$ from a pre-trained variational autoencoder – VAE (consists of an encoder $\mathcal{E}(\cdot)$ and decoder $\mathcal{D}(\cdot)$ in cascade) [40] to convert the input image $\mathbf{x}_{0}\in\mathbb{R}^{h\times w\times 3}$ to a latent image $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})\in\mathbb{R}^{\frac{h}{8}\times\frac{w}{8}\times d}$. Later, a time-conditional denoising UNet [70] $\mathcal{F}_{\theta}(\cdot)$ is trained to denoise directly on the latent space. SD’s UNet architecture consists of $12$ encoding layers, $1$ middle layer, and $12$ skip-connected decoding layers [69]. Among the $12$ encoding and $12$ decoding layers, there are $4$ downsampling $\{\mathcal{F}_{\mathbf{d}}^{1},\mathcal{F}_{\mathbf{d}}^{2},\mathcal{F}_{\mathbf{d}}^{3},\mathcal{F}_{\mathbf{d}}^{4}\}$ and $4$ upsampling $\{\mathcal{F}_{\mathbf{u}}^{1},\mathcal{F}_{\mathbf{u}}^{2},\mathcal{F}_{\mathbf{u}}^{3},\mathcal{F}_{\mathbf{u}}^{4}\}$ layers respectively. $\mathcal{F}_{\theta}(\cdot)$ takes three inputs – (i) the $t^{\text{th}}$ step noisy latent $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, (ii) the scalar time-step value $t$, and (iii) textual embedding $\mathbf{T}_{p}=\mathcal{T}(\mathbf{p})$in $\mathbb{R}^{77\times d}$. Both time-embedding and textual embedding influence the intermediate feature maps of the UNet via cross-attention [69]. $\mathcal{F}_{\theta}(\cdot)$ then predicts the noise map $\hat{\epsilon}_{t}$ (same size as $\mathbf{z}_{t}$) corresponding to that time-step. Mathematically,
$\mathcal{F}_{\theta}:(\mathbf{z}_{t},t,\mathbf{T}_{p})\rightarrow\hat{\epsilon}_{t}\vspace{-0.15cm}$ (1)
@@ 6 Background: Baseline Zero-Shot SBIR
$\mathcal{L}_{\mathrm{trip}}=\mathtt{max}\{0,\mu+\delta(f_{s},f_{p})-\delta(f_{s},f_{n})\}\vspace{-0.2cm}$ (2)
@@ 8.1 Result Analysis
Performance in Low-Data Scenario.  Existing ZS-(FG)-SBIR models are limited by the availability of paired sketch-photo training data [78, 103, 20, 16]. To test our method in a low-data scenario, we experiment by varying the number of training data ($10\%$, $30\%$, $50\%$, $70\%$, $100\%$) per class from Sketchy [81]. Our method remains relatively stable, outperforming (Fig. 4) other baseline and SOTAs significantly, preserving almost at par accuracy with full-scale training, by harnessing the out-of-distribution generalisation [71, 90, 100] capability of pre-trained SD [69].
@@ 10 Conclusion
Subhadeep Koley1,2    Ayan Kumar Bhunia1    Aneeshan Sain1,2    Pinaki Nath Chowdhury1,2
@@ C. Ablating Stable Diffusion versions
We ablate multiple SD [69] versions on Sketchy [81] dataset in Tab. 6. While SD v1.x models utilise CLIP [65] text encoder during their pre-training, v2.x models resort to OpenCLIP [9]. Evidently, SD v2.x models perform better than v1.x ones with v2.1 achieving highest scores. This is likely due to v2.x models’ adaptation of the much larger-scale OpenCLIP [9] encoder during pre-training.

### 20 Mar 2024 v2

@@ 3 Revisiting Text-to-Image Diffusion Models
Stable Diffusion Architecture.  Given an image and text-prompt pair ($\mathbf{x,p}$), SD [69] first uses the encoder $\mathcal{E}(\cdot)$ from a pre-trained variational autoencoder – VAE (consists of an encoder $\mathcal{E}(\cdot)$ and decoder $\mathcal{D}(\cdot)$ in cascade) [40] to convert the input image $\mathbf{x}_{0}\in\mathbb{R}^{h\times w\times 3}$ to a latent image $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})\in\mathbb{R}^{\frac{h}{8}\times\frac{w}{8}\times d}$. Later, a time-conditional denoising UNet [70] $\mathcal{F}_{\theta}(\cdot)$ is trained to denoise directly on the latent space. SD’s UNet architecture consists of $12$ encoding layers, $1$ middle layer, and $12$ skip-connected decoding layers [69]. Among the $12$ encoding and $12$ decoding layers, there are $4$ downsampling $\{\mathcal{F}_{\mathbf{d}}^{1},\mathcal{F}_{\mathbf{d}}^{2},\mathcal{F}_{\mathbf{d}}^{3},\mathcal{F}_{\mathbf{d}}^{4}\}$ and $4$ upsampling $\{\mathcal{F}_{\mathbf{u}}^{1},\mathcal{F}_{\mathbf{u}}^{2},\mathcal{F}_{\mathbf{u}}^{3},\mathcal{F}_{\mathbf{u}}^{4}\}$ layers respectively. $\mathcal{F}_{\theta}(\cdot)$ takes three inputs – (i) the $t^{\text{th}}$ step noisy latent $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, (ii) the scalar time-step value $t$, and (iii) textual embedding $\mathbf{T}_{p}=\mathcal{T}(\mathbf{p})$in $\mathbb{R}^{77\times d}$. Both time-embedding and textual embedding influence the intermediate feature maps of the UNet via cross-attention [69]. $\mathcal{F}_{\theta}(\cdot)$ then predicts the noise map $\hat{\epsilon}_{t}$ (same size as $\mathbf{z}_{t}$) corresponding to that time-step. Mathematically,
$\mathcal{F}_{\theta}:(\mathbf{z}_{t},t,\mathbf{T}_{p})\rightarrow\hat{\epsilon}_{t}\vspace{-0.1cm}$ (1)
@@ 6 Background: Baseline Zero-Shot SBIR
$\mathcal{L}_{\text{trip}}=\mathtt{max}\{0,\mu+\delta(f_{s},f_{p})-\delta(f_{s},f_{n})\}\vspace{-0.2cm}$ (2)
@@ 8.1 Result Analysis
Performance in Low-Data Scenario.  Existing ZS-(FG)-SBIR models are limited by the availability of paired sketch-photo training data [78, 103, 20, 16]. To test our method in a low-data scenario, we experiment by varying the number of training data ($10\%$, $30\%$, $50\%$, $70\%$, $100\%$) per class from Sketchy [81]. Our method remains relatively stable, outperforming (Fig. 4) other baseline and SoTAs significantly, preserving almost at par accuracy with full-scale training, by harnessing the out-of-distribution generalisation [71, 90, 100] capability of pre-trained SD [69] model.
@@ 10 Conclusion
Subhadeep Koley1,2    Ayan Kumar Bhunia1    Aneeshan Sain1    Pinaki Nath Chowdhury1
@@ C. Ablating Stable Diffusion versions
We ablate multiple SD [69] versions on Sketchy [81] dataset in Tab. 6. While SD v1.x models utilise CLIP [65] text encoder during their pre-training, v2.x models resort to much larger-scale OpenCLIP [9]. Evidently, SD v2.x models perform better than v1.x ones with v2.1 achieving the highest score. This is likely due to v2.x models’ adaptation of the much larger-scale OpenCLIP [9] encoder during pre-training.